OpenStudy (anonymous):

Let f(x) = 3x2 + x – 3 and g(x) = x2 – 5x + 1. Find f(x) – g(x). a. 2x2 – 4x – 2 b.2x2 – 4x – 4 c.2x2 + 6x – 2 d.2x2 + 6x – 4

OpenStudy (anonymous):

Is that x2 or x^2?

OpenStudy (anonymous):

Not that it matters. 3x2-x2=2x2 x-(-5x)=6x -3-(1)=-4

OpenStudy (anonymous):

Put them all together: 2x2+6x-4

OpenStudy (precal):